# Generating Physically Stable and Buildable LEGO® Designs from Text

### v1

@@ 6 Discussion and Limitations
In this work, we have introduced LegoGPT, an autoregressive model for generating LEGO designs from text prompts. Our method learns to predict the next brick sequentially while ensuring physical stability and buildability. We have shown that our method outperforms LLM backbone models and several recent text-to-3D generation methods.
Limitations.
Though our method outperforms existing methods, it still has several limitations. First, due to limited computational resources, we have not explored the largest 3D dataset. As a result, our method is restricted to generating designs within a $20\times 20\times 20$ grid across 21 categories. Future work includes scaling up model training on larger, more diverse datasets, such as Objaverse-XL [8] at higher grid resolutions. Training on large-scale datasets can also improve generalization to out-of-distribution text prompts.
Second, our method currently supports a fixed set of commonly used LEGO bricks. In future work, we plan to expand the brick library to include a broader range of dimensions and brick types, such as slopes and tiles, allowing for more diverse and intricate LEGO designs.
Acknowledgments.
We thank Minchen Li, Ken Goldberg, Nupur Kumari, Ruihan Gao, and Yihao Shi for their discussions and help. We also thank Jiaoyang Li, Philip Huang, and Shobhit Aggarwal for developing the bimanual robotic system. This work is partly supported by the Packard Foundation, Cisco Research Grant, and Amazon Faculty Award. This work is also in part supported by the Manufacturing Futures Institute, Carnegie Mellon University, through a grant from the Richard King Mellon Foundation. KD is supported by the Microsoft Research PhD Fellowship.
@@ Acknowledgments.
Appendix A LegoGPT Implementation Details
Captioning Details.
The complete prompt template used for GPT-4o caption generation is as follows:
@@ 5. Excludes medium-related terms (model, render, design)
StableText2Lego Details.
We generate StableText2Lego from the objects in ShapeNetCore [2]. While ShapeNetCore provides both voxel and mesh representations, we find that working directly with mesh data preserves better geometric details. We voxelize the 3D mesh into a $20\times 20\times 20$ grid representation and generate the LEGO structure using a split-and-remerge legolization algorithm. Split-and-remerge legolization has been studied in prior work [45]. But instead of initializing the voxel with $1\times 1$ unit bricks and randomly merging them into larger bricks [45], we initialize with heuristics by directly placing bricks to fill all the voxels, i.e., prioritizing placing 1) larger bricks and 2) bricks that connect multiple other bricks. After initializing the layout, we apply the new stability analysis [37] to identify weak parts in the structure and rearrange the local brick layout if necessary. In addition, to further improve data quality and diversity, we introduce randomness in the legolization process. Specifically, when initializing the LEGO structure, we randomly permute the order of brick placement within the same heuristic priority and generate two different structures for each object. With the mesh-to-LEGO and structure augmentation techniques, we generate $62{,}000+$ LEGO structures covering the $21$ categories with $31{,}218$ unique 3D objects from ShapeNetCore. Among the selected 3D objects, $\sim$92(i.e., $28{,}822$) of them have at least one stable LEGO design, which offer $47{,}000+$ stable LEGO layouts. Our StableText2Lego dataset significantly expands upon the previous StableLego dataset [37] in several key aspects: it contains >3×>3$\times$> 3 × more stable unique 3D objects with >5×>5$\times$> 5 × more stable LEGO structures compared to $8{,}000+$ in StableLego, spans a diverse set of 21 object categories, and provides detailed geometric descriptions for each shape.
Figure 9: Given a prompt and several example LEGO designs in text format, a pre-trained LLaMA model can generate LEGO designs with some structure.
Figure 10: Novelty analysis. For each structure generated by LegoGPT, we find the closest structure in the training dataset as measured by Chamfer distance in voxel space. The generated structures are distinct from their nearest neighbors, indicating low memorization.
Training details.
The full prompt that we use to construct our instruction fine-tuning dataset is as follows:
@@
Allowed brick dimensions are 2x4, 4x2, 2x6, 6x2, 1x2, 2x1, 1x4, 4x1, 1x6, 6x1, 1x8, 8x1, 1x1, 2x2. All bricks are 1 unit tall.
@@
Inference Details.
The sampling temperature is set to 0.6 for all experiments. During brick-by-brick rejection sampling, to mitigate repeated generation of rejected bricks, we increase the temperature by 0.01 each time the model generates a brick that has already been rejected.
To force each output brick to be in the format “{h}×{w} ({x},{y},{z})”, we sample only from the set of valid tokens during each step. For example, the first token must be a digit, the second an ‘x’, and so on. This has little to no effect on the output of LegoGPT, but helps force the baseline pre-trained models to output LEGO bricks in the specified format.
@@
Novelty Analysis.
For each generated structure, we find its closest structure in the training dataset, measured by computing the Chamfer distance in voxel space. As seen in Figure 10, the generated structures are distinct from their nearest neighbors in the dataset, confirming that our model can create novel designs rather than simply memorizing the training data.
In-context Learning.
We use LLaMA-3.2-1B-Instruct [11] as our base model, chosen for its computational efficiency. The in-context learning pipeline is shown in Figure 9, where the base model can generate LEGO-like designs through in-context learning, while suffering from collisions and disconnectivity. We did not use rejection sampling or rollback when evaluating zero-shot or few-shot generation, as doing so resulted in an excessive number of rejections and a sharp increase in generation time.
@@ Appendix B Robotic Assembly
We demonstrate automated assembly using a dual-robot-arm system as shown in Figure 8. The system consists of two Yaskawa GP4 robots, each equipped with an ATI force-torque sensor. A calibrated LEGO plate is placed between them, and the robots use the bricks initially placed on the plate to construct the LEGO structure. Given a $B=[b_{1},b_{2},\dots,b_{N}]$, we employ the action mask in [39] with assembly-by-disassembly search [71] to generate a physically executable assembly sequence for the robots, i.e., reordering the brick sequence so that 1) each intermediate structure is physically stable by itself, 2) each intermediate structure is stable under the robot operation, and 3) each assembly step is executable within the system’s dexterity. With the executable assembly sequence, an asynchronous planner [23] distributes the assembly tasks to the robots, plans the robots’ movements, and coordinates the bimanual system to construct the LEGO structure. The robots use the end-of-arm tool and the manipulation policy presented in [38, 23] with closed-loop force control to robustly manipulate LEGO bricks and construct the structure.
@@ Appendix C Manual Assembly
Figure 11: Manual Assembly. Demonstrations of manual assembly of generated LEGO structures.

### v2

@@ 6 Discussion and Limitations
In this work, we have introduced BrickGPT, an autoregressive model for generating interconnecting toy brick structures from text prompts. Our method learns to predict the next brick sequentially while ensuring physical stability and buildability. We have shown that our method outperforms LLM backbone models and several recent text-to-3D generation methods.
Limitations.   Though our method outperforms existing methods, it still has several limitations. First, due to limited computational resources, we have not explored the largest 3D dataset. As a result, our method is restricted to producing designs within a $20\times 20\times 20$ grid across 21 categories, while recent 3D generation methods can create a wider variety of objects. Future work includes scaling up model training at higher grid resolutions on larger, more diverse datasets, such as Objaverse-XL [7]. Training on large-scale datasets can also improve generalization to out-of-distribution text prompts.
Second, our method currently supports a fixed set of commonly used toy bricks. In future work, we plan to expand the brick library to include a broader range of dimensions and brick types, such as slopes and tiles, allowing for more diverse and intricate designs.
Acknowledgments.   We thank Minchen Li, Ken Goldberg, Nupur Kumari, Ruihan Gao, and Yihao Shi for their discussions and help. We also thank Jiaoyang Li, Philip Huang, and Shobhit Aggarwal for developing the bimanual robotic system. This work is partly supported by the Packard Foundation, Cisco Research Grant, and Amazon Faculty Award. This work is also in part supported by the Manufacturing Futures Institute, Carnegie Mellon University, through a grant from the Richard King Mellon Foundation. KD is supported by the Microsoft Research PhD Fellowship.
@@
Appendix A BrickGPT Implementation Details
Captioning Details.   The complete prompt template used for GPT-4o caption generation is as follows:
@@ 5. Excludes medium-related terms (model, render, design)
StableText2Brick Details.   We generate StableText2Brick starting from the objects in ShapeNetCore [2]. While ShapeNetCore provides both voxel and mesh representations, we find that working directly with mesh data better preserves geometric details. We voxelize the 3D mesh into a $20\times 20\times 20$ grid representation and generate its brick structure using a delete-and-rebuild algorithm. This algorithm is similar to that studied in prior work [46]. However, instead of initializing the structure by filling its voxels with $1\times 1$ unit bricks and randomly merging them into larger bricks, we greedily place bricks to fill the voxels layer-by-layer from bottom to top. We prioritize placing 1) bricks that are only partially supported by bricks on the layer below, 2) bricks that touch multiple bricks on the layer below, 3) large bricks, and 4) bricks of the opposite orientation from bricks on the layer below.
After initializing the structure, we apply the new stability analysis algorithm [38] to iteratively identify weak regions, delete their bricks, and rebuild them by greedily placing bricks, prioritizing 1) bricks that connect multiple disconnected components and 2) large bricks. This process is stochastic; we choose weak regions randomly as per [46], and we break ties between two bricks of the same heuristic value randomly. Using this delete-and-rebuild algorithm, we generate two different structures for each object.
With our delete-and-rebuild mesh-to-brick algorithm, we generate $62{,}000+$ brick structures covering the $21$ categories with $31{,}218$ unique 3D objects from ShapeNetCore. Among the selected 3D objects, $\sim$92(i.e., $28{,}822$) of them have at least one stable brick design, which offer $47{,}000+$ stable layouts. Our StableText2Brick dataset significantly expands upon the previous StableLego dataset [38] in several key aspects: it contains >3×>3$\times$> 3 × more stable unique 3D objects with >5×>5$\times$> 5 × more stable structures compared to $8{,}000+$ in StableLego, spans a diverse set of 21 object categories, and provides detailed geometric descriptions for each shape.
Figure 9: Given a prompt and several example structures in text format, a pre-trained LLaMA model can generate brick designs with some structure.
Figure 10: Novelty analysis. For each structure generated by BrickGPT, we find the closest structure in the training dataset as measured by Chamfer distance in voxel space. The generated structures are distinct from their nearest neighbors, indicating low memorization.
Training details.   The full prompt that we use to construct our instruction fine-tuning dataset is as follows:
@@
Allowed brick dimensions are $1\times 1$, $1\times 2$, $2\times 1$, $1\times 4$, $4\times 1$, $1\times 6$, $6\times 1$, $1\times 8$, $8\times 1$, $2\times 2$, $2\times 4$, $4\times 2$, $2\times 6$, $6\times 2$. All bricks are 1 unit tall.
@@
Inference Details.   The sampling temperature is set to 0.6 for all experiments. During brick-by-brick rejection sampling, to mitigate repeated generation of rejected bricks, we increase the temperature by 0.01 each time the model generates a brick that has already been rejected.
To force each output brick to be in the format “{h}×{w} ({x},{y},{z})”, we sample only from the set of valid tokens during each step. For example, the first token must be a digit, the second an ‘x’, and so on. This has little to no effect on the output of BrickGPT, but helps force the baseline pre-trained models to output bricks in the specified format.
@@
Novelty Analysis.   For each generated structure, we find its closest structure in the training dataset, measured by computing the Chamfer distance in voxel space. As seen in Figure 10, the generated structures are distinct from their nearest neighbors in the dataset, confirming that our model can create novel designs rather than simply memorizing the training data.
In-context Learning.   We use LLaMA-3.2-1B-Instruct [11] as our base model, chosen for its computational efficiency. The in-context learning pipeline is shown in Figure 9, where the base model can generate brick structures through in-context learning, while suffering from collisions and disconnectivity. We do not use rejection sampling or rollback when evaluating zero-shot or few-shot generation, as doing so results in an excessive number of rejections and a sharp increase in generation time.
@@ Appendix B Robotic Assembly
We demonstrate automated assembly using a dual-robot-arm system as shown in Figure 8. The system consists of two Yaskawa GP4 robots, each equipped with an ATI force-torque sensor. A calibrated baseplate is placed between them, and the robots use the bricks initially placed on the plate to construct the brick structure. Given a brick structure $B=[b_{1},b_{2},\dots,b_{N}]$, we employ the action mask in [40] with assembly-by-disassembly search [73] to generate a physically executable assembly sequence for the robots, i.e., reordering the brick sequence so that 1) each intermediate structure is physically stable by itself, 2) each intermediate structure is stable under the robot operation, and 3) each assembly step is executable within the system’s dexterity. With the executable assembly sequence, an asynchronous planner [24] distributes the assembly tasks to the robots, plans the robots’ movements, and coordinates the bimanual system to construct the brick structure. The robots use the end-of-arm tool and the manipulation policy presented in [39, 24] with closed-loop force control to robustly manipulate bricks and construct the structure.
@@ Appendix C Manual Assembly
Figure 11: Manual Assembly. We demonstrate manual assembly of generated structures using LEGO bricks.